# Searching for Exotrojans in Pulsar Binary Systems

### v1

@@ 3.2.2 Libration Priors
We adopt uninformative priors on the libration parameters, $\mathcal{T}$, $\nu$, $\theta$, and $\varphi$. We also adopt the same priors on the red noise as in (L. Dey et al., 2025). We do not search for $n_{\rm b}$ in our MCMC approach as the mean motion is already well constrained by the timing model.
@@
  • We choose a log-uniform distribution for $\mathcal{T}$ of $\log{(\mathcal{T}/\rm{s})}\sim\mathcal{U}(-12,\,-4)$. The lower limit of 12-12 was chosen because such a signal strength is far beneath the noise floor of NG15yr MSPs, which is on the order of 0.1–1 microseconds (G. Agazie et al., 2023). The 4-4 upper limit is in place because a signal of such strength would already easily be detected by eye in the TOAs residuals.
@@
  • Our prior on $\nu$ is $\nu\sim\mathcal{U}(\nu_{\rm min},\,\nu_{\rm max})$. The maximum real-valued $\nu$ in Equation 11 is $n_{\rm b}/\sqrt{2}$. Therefore, this is our $\nu_{\rm max}$. $\nu_{\rm min}$ is determined by finding the smallest $\mu$ for a given binary mass function (Equation 17) and inserting that $\mu$ into Equation 10. Such a $\mu$ is found by setting $\sin{I}=1$ and the pulsar mass to the most massive discovered pulsar ($\sim 2.5\,\mathrm{M}_{\odot}$; N. Clifford 2019).
@@
  • We choose the prior $\theta\sim\mathcal{U}(\pi/3,\,2\pi/3)$ because these bound the physical range of $\theta$ according to Equation 9, given $\delta\in[0,\,1]$. We also investigate the $\theta\sim\mathcal{U}(0,\,2\pi)$ prior to account for any nonphysical signals that may enter our physically-motivated prior space due to uncertainties in $\theta$. Lastly, we look in $\log{\xi}$ space, where $\theta=\pi/3+\xi$, with a $\log{\xi}\sim\mathcal{U}(-6.2,\,\log{(\pi/3)})$ prior. The minimum of 6.2-6.2 is chosen because this gives the corresponding $\theta$ when the Trojan is 1 lunar mass and the dominant companion is $0.05\,\rm{M}_{\odot}$, the most massive companion satisfying Equation 1 when $m_{\rm p}=1.35\,\rm{M}_{\odot}$. A search in $\log{\xi}$ allows us to probe small Trojan masses in the $L_{5}$ Lagrange point (Section 4.3). We do not carry out an analogous $\theta=2\pi/3-\xi^{\prime}$ search over a uniform $\log{\xi^{\prime}}$ prior because none of our positive detections are near $\theta=2\pi/3$.
@@
4 Results
4.1 PSR 1641+8049 Superior Conjunction Inconsistency Result
@@
Using the optical lightcurves of the PSR J1641+8049 system, we identify two times of superior conjunction corresponding to peak light. These are $T_{\rm o,s}=\{58640.0974(15)\,,58640.1883(15)\}$ MJD. Note the binary period is $0.09087395384(1)$ day. The in-system Romer delay is weaker than the uncertainties, so no correction is needed for $T_{\rm o,s}$. The corresponding times of superior conjunction as predicted by radio pulsar timing are $T_{\rm r,s}=\{58640.09736,\,58640.18823\}$ MJD, where we omit unnecessary additional digits of accuracy. Here, we use the $T_{\rm r,s}=T_{\rm asc}-P_{\rm b}/4$ equation and subtracted integer multiples of $P_{\rm b}$ so that $T_{\rm r,s}$ is as close as possible to $T_{\rm o,s}$. The results are visually summarized in Figure 3.
@@
These different times of superior conjunction give an average time difference of $\Delta t=4.6\pm 94$ s, which corresponds to $\Delta m=66\pm 1400$M through Equation 2. Thus, we constrain the mass difference to be no higher than 8 MJ between the triangular Lagrange points at the 95% confidence level. Comparing times of superior conjunction also confirms, within the uncertainties, that the companion of PSR J1641+8049 is being heated symmetrically (see D. Kandel & R. W. Romani 2020; G. Voisin et al. 2020 for discussions on asymmetric black widow companion heating). Combining both a non-zero $\Delta m$ and asymmetric heating in such a way that their effects on the time of superior conjunction cancel is also possible.
Figure 3: HiPERCAM light curve data of the PSR J1641+8049 system for each Sloan band. The data was first presented in D. Mata Sánchez et al. (2023). We show the best-fit model for the $r^{\prime}$ band as an example, with the remaining model curves shown in A. Y. Kirichenko et al. (2024). Every band except $u^{\prime}$ was used in calculating $T_{\rm o,s}$. The vertical dashed line marks the time of superior conjunction as predicted by radio timing. The vertical cyan band encloses the 68% confidence interval of the time of superior conjunction as calculated from the optical light curve.
4.2 Librations Results
@@
With the exceptions of PSRs J0023+0923 and J1705-1903, we find zero evidence of exotrojans in the NG15yr pulsar systems analyzed. In Figure 4 we showcase a corner plot of the posteriors for PSR J0610-2100, which is characteristic of the other five clear non-detections. To emphasize that these are indeed non-detections, we calculate the Bayes factor, $\mathcal{B_{T}}$, for each NG15yr pulsar. We use the Savage-Dickey formalism (J. M. Dickey, 1971) adapted for our model,
@@
Upon a positive detection of the signal given by Equation 3.2, we employ another Bayes factor,
@@
Figure 4: The posterior space of the PSR J0610-2100 libration analysis. Each histogram represents the marginalized posterior of the relevant parameter. Each contour plot represents the joint-marginalized posterior of the two relevant parameters. This is a clear non-detection with almost flat posteriors across all parameters. There is a slight preference for $\log{(\mathcal{T}/\rm{s})}\sim 10^{-6}$, which is similar to the noise floor of this pulsar (G. Agazie et al., 2023), though $\log{(\mathcal{T}/\rm{s})}$ is agnostic to the other three parameters. Furthermore, the Bayes factor for a non-zero $\log{(\mathcal{T}/\rm{s})}$ is less than 1, so no librations ($\mathcal{T}\equiv 0$) are probabilistically preferred.
@@
[…]
@@
Due to mostly non-detections, we place upper limits (ULs) on the libration strength ($\mathcal{T_{\rm UL}}$) and make additional assumptions on the pulsar and pulsar Trojan populations to map the marginalized posterior on $\mathcal{T}$ to a Trojan mass UL ($m_{\rm T,UL}$; see Appendix B for details on the $\mathcal{T}\mapsto m_{\rm T,UL}$ mapping). We choose our ULs to represent 95% credibility. As in L. Dey et al. (2025), we reweight the posterior on $\mathcal{T}$ to match the posterior expected when instead choosing a uniform, rather than log-uniform, prior on $\mathcal{T}$. Each $\mathcal{T_{\rm UL}}$ and $m_{\rm T,UL}$ is listed in the third and fourth columns, respectively, of Table 1.
@@
Despite positive detections of Equation 3.2 in the PSR J0023+0923 and J1705-1903 systems, we also place upper limits on them. In doing so, we assume the signals are not from true librations, and therefore, we follow the same procedure as for the other pulsars. Thus, the Trojan mass upper limits on these two systems are inconsistent with the Trojan masses as calculated in Table 2 because the upper limits assume a certain distribution in $z$, while the results in Table 2 can determine $z$ from the libration parameters. See Section 4.3 and Appendix B for more details.
4.3 PSRs J0023+0923 and J1705-1903 Libration Results
@@
We find weak evidence of exotrojans in the PSR J0023+0923 and J1705-1903 systems and showcase their posterior corner plots in Figures 5 and 6. For PSR J0023+0923, the posterior space is clearly multimodal, with evidence for sidereal-day-aliasing in $\nu$. Aliasing is an observational cadence effect due to the Nyquist-Shannon sampling theorem, where spurious frequencies can fit the data somewhat well in place of the true frequency. In our case, these different frequencies vary from the true frequency by integer multiples of $f_{\rm sidereal}\approx 1\,\rm{day}^{-1}$, because the NG15yr TOAs of PSR J0023+0923 were measured by the Arecibo Observatory, a transit instrument. Therefore, every TOA of PSR J0023+0923 was observed at roughly the same sidereal time. As mentioned in Section 2.2, the Nyquist-Shannon sampling theorem is weakened by the vast number of observations occurring at different librational phases, which is why one frequency, $\sim 3.82\,\rm{day}^{-1}$, stands out above the rest. PSR J1705-1903 only shows weak aliasing, because it was observed with the Green Bank Telescope. Another feature of both posterior spaces is that $\theta$ tends towards $\pi/3$, which corresponds to $\delta=0$. Therefore, any Trojans in these two systems will be in the $L_{5}$ Lagrange point, and thus $m_{\rm c}$ and $m_{\rm T}$ should be swapped in Figure 2.
@@
Figure 5: The posterior space of the PSR J0023+0923 libration analysis. Each histogram represents the marginalized posterior of the relevant parameter. Each contour plot represents the joint-marginalized posterior of the two relevant parameters. With a Bayes factor of 11.5(2), we have a tentative detection of the signal in Equation 3.2. However, we find that the origin of this signal is unlikely to be caused by a Trojan due to a nonphysical Gaussian-average value for $\theta$. See the text for more details.
@@
In Figure 8 we show the PSR J0023+0923 posterior space in the $\pm$ 0.05% neighborhood of $\nu=3.81759\,\rm{day}^{-1}$. We find that the $\theta$ posterior is fully consistent with that of a Gaussian tail, where the Gaussian is centered at $\theta\approx 0.64$, implying $\delta\approx-0.66$. Furthermore, we find a Gaussian in $\theta$ centered at the same value when searching for a Trojan signal when using a $\theta\sim\mathcal{U}(0,\,2\pi)$ prior, rather than the physically valid $\theta\sim\mathcal{U}(\pi/3,\,2\pi/3)$ prior. We also search in $\log{\xi}$ space (see Section 3.2.2) and find no preferred $\xi$ (see Figure 7). In fact, we find $\mathcal{B}_{\xi}<1$ for both pulsars. We therefore conclude that the signals associated with the $\nu\approx 3.82\,\rm{day}^{-1}$ and $3.02\,\rm{day}^{-1}$ libration frequencies for PSR J0023+0923 and J1705-1903, respectively, are unlikely to be caused by Trojans. However, for instructive purposes, we carry out the following calculations as if the signals were from a Trojan in each system.
Figure 7: The same as Figure 5 except $\log{\xi}$ is searched over uniformly.
In this paragraph, we only use samples where $\nu$ is in the $\pm$ 0.05% neighborhood of $\nu=3.81759\,\rm{day}^{-1}$ for PSR J0023+0923 and $3.01600\,\rm{day}^{-1}$ for PSR J1705-1903. Using Equation 12 on the $\theta$ samples and Equations 11 and 15 on the $\nu$ samples, we find $\delta$ and $q\equiv m_{\rm p}/m_{\rm c}$, placing them in the second and third columns, respectively, of Table 2. We then use Equation 14 to calculate $z$, which we include in the fourth column of Table 2. Swapping $m_{\rm c}$ and $m_{\rm T}$ in the definition of $\delta$ so that $m_{\rm c}$ is the more massive body, and assuming the pulsar mass function of J. Antoniadis et al. (2016) (as in Appendix B), we find $m_{\rm T}$, $m_{\rm c}$, and $I$, placing them in the remaining columns of Table 2. Using an optical light curve analysis of the PSR J0023+0923 system, D. Mata Sánchez et al. (2023) find $q=59_{-11}^{+10}$ and $I=42_{-3}^{+4}$°. The inconsistency of their measurements with ours is another indication that our detected signal is not from a Trojan.
@@
Table 2: Libration parameters for PSRs J0023+0923 and J1705-1903 if their detected signals are from individual Trojans. Values are derived by swapping $m_{\rm c}$ and $m_{\rm T}$ in the definition of $\delta$ and assuming the pulsar mass function of J. Antoniadis et al. (2016).
@@
Figure 8: The same as Figure 5 except the posterior is restricted to the $\pm$ 0.05% neighborhood of $\nu=3.81759\,\rm{day}^{-1}$. The $\pm$ 0.05% width was chosen so that several standard deviations were included. A Fisher information analysis reveals this width to be the expected uncertainty in $\nu$. We define $\nu^{\prime}=\nu-3.81759\,\rm{day}^{-1}$. As stated in the text, we find it unlikely that this signal is due to a Trojan.
4.4 Librating Trojan Prevalence

### v2

@@ II.2.2 Libration Priors
We adopt uninformative priors on the libration parameters, $\mathcal{T}$, $\nu$, $\theta$, and $\varphi$. We also adopt the same priors on the red noise as in (Dey et al., 2026). We do not search for $n_{\rm b}$ in our MCMC approach as the mean motion is already well constrained by the timing model.
@@
  • We choose a log-uniform distribution for $\mathcal{T}$ of $\log{(\mathcal{T}/\rm{s})}\sim\mathcal{U}(-12,\,-4)$. The lower limit of 12-12 was chosen because such a signal strength is far beneath the noise floor of NG15yr MSPs, which is on the order of 0.1–1 microseconds (Agazie et al., 2023). The 4-4 upper limit is in place because a signal of such strength would already easily be detected by eye in the TOA residuals.
@@
  • Our prior on $\nu$ is $\nu\sim\mathcal{U}(\nu_{\rm min},\,\nu_{\rm max})$. The maximum real-valued $\nu$ in Equation 15 is $n_{\rm b}/\sqrt{2}$. Therefore, this is our $\nu_{\rm max}$. The minimum value, $\nu_{\rm min}$, is determined by finding the smallest $\mu$ for a given binary mass function (Equation 28) and inserting that $\mu$ into Equation 13. Such a $\mu$ is found by setting $\sin{I}=1$ and the pulsar mass to the most massive discovered pulsar ($\sim 2.5\,\mathrm{M}_{\odot}$; Clifford 2019).
@@
  • We choose the prior $\theta\sim\mathcal{U}(\pi/3,\,2\pi/3)$ because these bound the physical range of $\theta$ according to Equation 17, given $\delta\in[0,\,1]$. We also investigate the $\theta\sim\mathcal{U}(0,\,2\pi)$ prior to account for any nonphysical signals that may enter our physically-motivated prior space due to uncertainties in $\theta$. Lastly, we look in $\log{\xi}$ space, where $\theta=\pi/3+\xi$, with a $\log{\xi}\sim\mathcal{U}(-6.2,\,\log{(\pi/3)})$ prior. The minimum of 6.2-6.2 is chosen because this gives the corresponding $\theta$ when the Trojan is 1 lunar mass and the secondary is $0.05\,\rm{M}_{\odot}$, the most massive companion satisfying Equation 1 when $m_{\rm p}=1.35\,\rm{M}_{\odot}$. A search in $\log{\xi}$, where
    $\xi=\theta-\pi/3,$ (29)
    allows us to probe small Trojan masses in the $L_{5}$ Lagrange point (Section IV.3). We do not carry out an analogous $\theta=2\pi/3-\xi^{\prime}$ search over a uniform $\log{\xi^{\prime}}$ prior because none of our positive detections are near $\theta=2\pi/3$.
@@
III Selected Pulsars and their Observational Data
According to the Australia Telescope National Facility Pulsar Catalog (Manchester et al., 2005), [^3] https://www.atnf.csiro.au/research/pulsar/psrcat there are 72 pulsar binary systems that satisfy Equation 1 when $m_{\rm T}=0$, $m_{\rm p}=1.35\,$M, and $m_{\rm c}$ is determined by the Newtonian mass function (Equation 28) when $\sin{I}=1$, where $I$ is the pulsar system’s orbital inclination with respect to the plane of the sky. We define $m_{\rm c,\,\min}$ as the minimum secondary mass when $m_{\rm p}=1.35\,$M. Note that $m_{\rm c,\,\min}=m_{\rm c}$ when $I=90^{\circ}$. Choosing $m_{\rm T}=0$ in evaluating Equation 1 gives us systems that may have Trojans, provided they are not too massive. Any discovered pulsar Trojans would have non-zero masses that still satisfy Equation 1. Out of these 72 systems, we restrict this study to PSR J1641+8049 and the eight systems that are in the NG15yr dataset (see Table 1) due to their high observation cadence, long timing baseline, and availability of the dataset. Other datasets also satisfy these criteria (e.g., Antoniadis et al., 2023; Zic et al., 2023), but we leave those analyses for future work due to differences in the noise modeling and observatories. Useful properties of these pulsars can be found in Table 1. For PSR J1641+8049, we use the method developed by Ford and Gaudi (2006) to seek out Trojan-caused inconsistencies between the pulsar-companion conjunction times as measured by optical lightcurves of the system (Kirichenko et al., 2024) and those predicted by radio pulsar timing (see Section II.1). For the eight NG15yr pulsars of interest, we adapt the method of Leleu et al. (2015) to look for TOA residuals caused by Trojan librations—a method that only requires TOAs (see Section II.2).
[…]
J1641+8049 2.1809748921(3) 0.0640753(2) $2.2(4)\times 10^{-5}$ 2.8 1.88 0.0404 GTC, CHIME
J0023+0923 3.3311794144(4) 0.03484142(4) $7(2)\times 10^{-6}$ 9.0 6.02 0.0164 AO
J0610-2100 6.864384046(3) 0.0734890(2) $1.8(8)\times 10^{-5}$ 3.4 9.41 0.0214 GBT
J0636+5128 1.5972322096(4) 0.00898636(5) $1.7(7)\times 10^{-5}$ 6.3 6.04 0.0069 GBT
J1705-1903 4.414896263(2) 0.1043598(1) $3(1)\times 10^{-6}$ 3.7 3.07 0.0412 GBT
J1719-1438 2.17695076(2) 0.0018216(4) $5(2)\times 10^{-4}$ 3.4 15.80 0.0011 GBT
J1745+1017 17.52579411(2) 0.0881697(6) $2(1)\times 10^{-5}$ 4.5 14.46 0.0137 AO
J2214+3000 9.999190702(3) 0.0590815(2) $2.1(6)\times 10^{-5}$ 8.4 7.31 0.0133 AO
J2234+0944 10.071840895(2) 0.0684298(1) $7(3)\times 10^{-6}$ 7.1 5.76 0.0154 AO
III.1 Optical Data
The first optical observations of the PSR J1641+8049 counterpart were reported by Lynch et al. (2018), revealing strong brightness variations tied to the binary period. The detection triggered further optical studies: using the High PERformance CAMera (HiPERCAM) installed at the 10.4m Gran Telescopio Canarias (GTC), Mata Sánchez et al. (2023) performed phase-resolved multi-band photometric observations and analysed the light curves of the system, yielding its binary parameters and the distance. Kirichenko et al. (2024) carried out independent phase-resolved multi-band optical studies with the Optical System for Imaging and low Resolution Integrated Spectroscopy (OSIRIS) at the GTC and performed a comparative light-curve analysis using both the HiPERCAM and OSIRIS datasets.
As demonstrated by both studies, the PSR J1641+8049 system light curves show a single-peaked sinusoidal-like modulation per orbital period with a maximum when the side of the companion heated by the neutron star is directed toward the observer (superior conjunction) and a minimum when the cold back side is visible (inferior conjunction). The system parameters obtained by Kirichenko et al. (2024) through the light curve analysis, including the distance to the source 4.6–4.8 kpc and the orbital inclination 56–59, are found to be consistent with those of Mata Sánchez et al. (2023). However, the optical flux of the source at the maximum brightness phase faded by a factor of $\sim 2$ as compared with the values reported in the Mata Sánchez et al. (2023) study. Despite this variability, the companion’s irradiated hemisphere remains among the hottest (8000–9500 K) observed in the known black widow systems.
III.2 Radio Pulsar Timing Data
The times of arrival for PSR J1641+8049 came from the CHIME/Pulsar backend (Amiri et al., 2021) on the Canadian Hydrogen Intensity Mapping Experiment (CHIME) radio telescope (Amiri et al., 2022). These were used in calculating the times of superior conjunction.
Of the eight selected NG15yr pulsars, four were observed with the Arecibo Observatory (AO; PSRs J0023+0923, J1745+1017, J2214+3000, and J2234+0944) and four with the Green Bank Telescope (GBT; PSRs J0610-2100, J0636+5128, J1705-1903, and J1719-1438). The TOAs of our selected pulsars cover time spans ($T_{\rm span}$) ranging from 3.4 yr (PSR J0610-2100) to 9.0 yr (PSR J0023+0923). The near-random telescope cadences with respect to libration phase (see Equations 12 and II.2), combined with the years-long timing baselines, ensure that the TOAs effectively sample all libration phases. Therefore, the Nyquist–Shannon sampling theorem criterion is relaxed, and the libration frequencies that are physically allowed (Section II.2.2) can be well-constrained. We show the TOA residual time series for PSRs J0023+0923 and J1705-1903 in Figure 4 as characteristic examples of residual time series. Residuals for the other pulsars can be found in Appendix C of Agazie et al. (2023). More details about the NG15yr dataset can be found in Agazie et al. (2023).
Figure 4: The pulse TOA residuals for PSRs J0023+0923 and J1705-1903. The error bars indicate the 68% uncertainty. These are the residuals after the ordinary timing model, but not red noise, has been accounted for (Section II.2.1). These two pulsars stand out above the others because their TOAs contain minor evidence for the Trojan libration signal (see Section IV.2).
IV Results
IV.1 PSR J1641+8049 Superior Conjunction Inconsistency Result
@@
Using the optical lightcurves of the PSR J1641+8049 system, we identify two times of superior conjunction corresponding to peak light. These are $T_{\rm o,s}=\{58640.0974(15)\,,58640.1883(15)\}$ MJD. Note the binary period is $0.09087395384(1)$ day. The in-system Romer delay is smaller than the uncertainties, so no correction is needed for $T_{\rm o,s}$. The corresponding times of superior conjunction as predicted by radio pulsar timing are $T_{\rm r,s}=\{58640.09735997(5),\,58640.18823392(5)\}$ MJD. Here, we use the $T_{\rm r,s}=T_{\rm asc}-P_{\rm b}/4$ equation and subtracted integer multiples of $P_{\rm b}$ so that $T_{\rm r,s}$ is as close as possible to $T_{\rm o,s}$. The results are visually summarized in Figure 1.
@@
We find no statistically significant difference in the TSCs, and constrain $|\Delta t|$ to be no higher than 190 s (95% confidence). Therefore, through Equation 2, $|\Delta m_{\rm T}|$ can no higher than 8 MJ between the triangular Lagrange points. Comparing times of superior conjunction also confirms, within the uncertainties, that the companion of PSR J1641+8049 is being heated symmetrically (see Kandel and Romani 2020; Voisin et al. 2020 for discussions on asymmetric black widow companion heating). Combining both a non-zero $\Delta m$ and asymmetric heating in such a way that their effects on the time of superior conjunction cancel is also possible.
IV.2 Librations Results
@@
With the exceptions of PSRs J0023+0923 and J1705-1903, we find zero evidence of exotrojans in the NG15yr pulsar systems analyzed. In Figure 5 we showcase a corner plot of the posteriors for PSR J0610-2100, which is characteristic of the other five clear non-detections. To emphasize that these are indeed non-detections, we calculate the Bayes factor, $\mathcal{B_{T}}$, for each NG15yr pulsar. We use the Savage-Dickey formalism (Dickey, 1971) adapted for our model,
@@
Upon a positive detection of the signal given by Equation II.2, we employ another Bayes factor,
@@
Figure 5: The posterior space of the PSR J0610-2100 libration analysis. Each histogram represents the marginalized posterior of the relevant parameter. Each contour plot represents the joint-marginalized posterior of the two relevant parameters. This is a clear non-detection with almost flat posteriors across all parameters. There is a slight preference for $\log{(\mathcal{T}/\rm{s})}\sim 10^{-6}$, which is similar to the noise floor of this pulsar (Agazie et al., 2023), though $\log{(\mathcal{T}/\rm{s})}$ is agnostic to the other three parameters. Furthermore, the Bayes factor for a non-zero $\log{(\mathcal{T}/\rm{s})}$ is less than 1, so no librations ($\mathcal{T}\equiv 0$) are probabilistically preferred.
@@
Table 2: Values of $\mathcal{T}_{\rm UL}$, Bayes factor $\mathcal{B_{T}}$ (with 68% uncertainty), $m_{\mathrm{T,UL}}$, and $\mathcal{B}_{\xi}$ for each pulsar. The uncertainty on the Bayes factor is due to the method in calculating the Bayes factor with non-infinite posterior chains. The ULs represent a 95% credibility. Every NG15yr pulsar except PSRs J0023+0923 and J1705-1903 has $\mathcal{B_{T}}<1$. We do not calculate $\mathcal{B}_{\xi}$ for these six pulsars. The upper limits placed on PSRs J0023+0923 and J1705-1903 were calculated using the same routine as the other pulsars and are independent of the analysis used in investigating a potential detection in the system in Section IV.3.
@@
Due to mostly non-detections, we place upper limits (ULs) on the libration strength ($\mathcal{T_{\rm UL}}$) and make additional assumptions on the pulsar and pulsar Trojan populations to map the marginalized posterior on $\mathcal{T}$ to a Trojan mass UL ($m_{\rm T,UL}$; see Appendix B for details on the $\mathcal{T}\mapsto m_{\rm T,UL}$ mapping). We choose our ULs to represent 95% credibility. Following Dey et al. (2026) we put $\mathcal{T}$ on a linear scale in order to correctly place an upper limit. This requires that we weight the $\mathcal{T}$ histogram bins by a factor of $\mathcal{T}$ to account for the Jacobian in converting between the $\log\mathcal{T}$ and $\mathcal{T}$ probability distributions. Hourihane et al. (2023) show that using a linear scale for $\mathcal{T}$ and weighting in this fashion ensure the upper limits do not depend on the $\mathcal{T}$ prior lower-bound. We place the values of $\mathcal{T_{\rm UL}}$ and $m_{\rm T}$ for each pulsar in third and fourth columns, respectively, of Table 2.
@@
Despite positive detections of Equation II.2 in the PSR J0023+0923 and J1705-1903 systems, we also place upper limits on them due to the fact that we suspect these are false-positives (see the following subsection). In doing so, we assume the signals are not from true librations, and therefore, we follow the same procedure as for the other pulsars. Thus, the Trojan mass upper limits on these two systems are inconsistent with the Trojan masses as calculated in Table 3 because the upper limits assume a certain distribution in $z$, while the results in Table 3 can determine $z$ from the libration parameters. See Section IV.3 and Appendix B for more details.
IV.3 PSRs J0023+0923 and J1705-1903 Libration Results
@@
We find weak evidence of exotrojans in the PSR J0023+0923 and J1705-1903 systems and showcase their posterior corner plots in Figures 6 and 7. For PSR J0023+0923, the posterior space is clearly multimodal, with evidence for sidereal-day-aliasing in $\nu$. Aliasing is an observational cadence effect due to the Nyquist-Shannon sampling theorem, where spurious frequencies can fit the data somewhat well in place of the true frequency. In our case, these different frequencies vary from the true frequency by integer multiples of $f_{\rm sidereal}\approx 1\,\rm{day}^{-1}$, because the NG15yr TOAs of PSR J0023+0923 were measured by the Arecibo Observatory, a transit instrument. As for all transit instruments, the Arecibo Observatory had limited pointing capabilities and required the sky to rotate over its field of view. Therefore, every TOA of PSR J0023+0923 was observed at roughly the same sidereal time. As mentioned in Section III.2, the Nyquist-Shannon sampling theorem is weakened by the vast number of observations occurring at different librational phases, which is why one frequency, $\sim 3.82\,\rm{day}^{-1}$, stands out above the rest. PSR J1705-1903 only shows weak aliasing, because it was observed with the Green Bank Telescope, which has full pointing capabilities. Another feature of both posterior spaces is that $\theta$ tends towards $\pi/3$, which corresponds to $\delta=0$. Therefore, any Trojans in these two systems will be in the $L_{5}$ Lagrange point, and thus $m_{\rm c}$ and $m_{\rm T}$ should be exchanged in Figure 2.
@@
Figure 6: The same as Figure 5 except for PSR J0023+0923. With a Bayes factor of 11.5(2), we have a tentative detection of the signal in Equation II.2. However, we find that the origin of this signal is unlikely to be caused by a Trojan due to a nonphysical Gaussian-average value for $\theta$. See the text for more details.
@@
In Figure 8 we show the PSR J0023+0923 posterior space in the $\pm$ 0.05% neighborhood of $\nu=3.81759\,\rm{day}^{-1}$. We find that the $\theta$ posterior is fully consistent with that of a Gaussian tail, where the Gaussian is centered at $\theta\approx 0.64$, implying $\delta\approx-0.66$. Furthermore, we find a Gaussian in $\theta$ centered at the same value when searching for a Trojan signal when using a $\theta\sim\mathcal{U}(0,\,2\pi)$ prior, rather than the physically valid $\theta\sim\mathcal{U}(\pi/3,\,2\pi/3)$ prior. We included a search over this non-physical prior in order to confirm that the source of the posterior is in the non-physical parameter space. We also search in $\log{\xi}$ space (see Section II.2.2) and find no preferred $\xi$ (see Figure 9). In fact, we find $\mathcal{B}_{\xi}<1$ for both pulsars. We therefore conclude that the signals associated with the $\nu\approx 3.82\,\rm{day}^{-1}$ and $3.02\,\rm{day}^{-1}$ libration frequencies for PSR J0023+0923 and J1705-1903, respectively, are unlikely to be caused by Trojans (see Section V for a discussion on possible other causes). However, for instructive purposes, we carry out the following calculations as if the signals were from a Trojan in each system.
Figure 8: The same as Figure 6 (PSR J0023+0923) except the posterior is restricted to the $\pm$ 0.05% neighborhood of $\nu=3.81759\,\rm{day}^{-1}$. The $\pm 0.05$% width was chosen so that several standard deviations were included. A Fisher information analysis reveals this width to be the expected uncertainty in $\nu$. We define $\nu^{\prime}=\nu-3.81759\,\rm{day}^{-1}$. As stated in the text, we find it unlikely that this signal is due to a Trojan.
Figure 9: The same as Figure 6 (PSR J0023+0923) except $\log{\xi}$ is searched over uniformly.
In this paragraph, we only use samples where $\nu$ is in the $\pm$ 0.05% neighborhood of $\nu=3.81759\,\rm{day}^{-1}$ for PSR J0023+0923 and $3.01600\,\rm{day}^{-1}$ for PSR J1705-1903. Using Equation 23 on the $\theta$ samples and Equations 15 and 26 on the $\nu$ samples, we find $\delta$ and $q\equiv m_{\rm p}/m_{\rm c}$, placing them in the second and third columns, respectively, of Table 3. We then use Equation 25 to calculate $z$, which we include in the fourth column of Table 3. Exchanging $m_{\rm c}$ and $m_{\rm T}$ in the definition of $\delta$ so that $m_{\rm c}$ is the more massive body, and assuming the pulsar mass function of Antoniadis et al. (2016) (as in Appendix B), we find $m_{\rm T}$, $m_{\rm c}$, and $I$, placing them in the remaining columns of Table 3. Using an optical light curve analysis of the PSR J0023+0923 system, Mata Sánchez et al. (2023) find $q=59_{-11}^{+10}$ and $I=42_{-3}^{+4}$°. The inconsistency of their measurements with ours is another indication that our detected signal is not from a Trojan.
@@
Table 3: Libration parameters for PSRs J0023+0923 and J1705-1903 if their detected signals are from individual Trojans. Values are derived by exchanging $m_{\rm c}$ and $m_{\rm T}$ in the definition of $\delta$ and assuming the pulsar mass function of Antoniadis et al. (2016).
@@
IV.4 Librating Trojan Prevalence